# Explainable deep learning framework for cancer therapeutic target prioritization leveraging PPI centrality and node embeddings

### v1

@@ 1 Introduction
The identification of essential genes and therapeutic targets represents a cornerstone of modern biomedical research, with profound implications for understanding fundamental biological processes and developing novel cancer treatments [Zhang2012_22848443]. Essential genes, defined as those critical for cellular survival and reproduction, have been extensively studied across various organisms, from bacteria to humans [Campos2019_31312416]. Contemporary advances increasingly emphasize human-based experimental and computational models as more predictive and ethically preferable alternatives to traditional animal systems. In cancer research, the prioritization of therapeutic targets is particularly crucial, as it directly influences drug discovery pipelines and treatment efficacy [Bazaga2020_32612205]. However, the complexity of biological systems and the vast diversity of human molecular and network data, coupled with population-level variability in genetic and epigenetic susceptibility, presents both a challenge and an opportunity for building integrative human-relevant models. [Safadi2024_38649399].
@@
Protein-protein interaction (PPI) networks have emerged as powerful tools for understanding gene essentiality and identifying potential therapeutic targets [Zhang2025_40691502]. These networks capture the intricate web of molecular interactions within cells, providing insights into cellular functions and disease mechanisms [Dai2020_32023848]. Centrality measures derived from PPI networks, such as degree, betweenness, and closeness centrality, have been shown to correlate with gene essentiality [Fang2018_29958434]. Moreover, advancements in network embedding techniques, such as Node2Vec, have enabled the extraction of latent topological features that enhance predictive accuracy [Dai2020_32023848].
@@
Previous prediction approaches often relied on isolated data modalities or species-agnostic features, lacking the combinatorial integration across multiple in silico modalities that can capture the biological complexity of human systems [Zhang2020_32936825]. For instance, DeepHE, a deep learning-based method, integrates sequence features with PPI network data to predict human essential genes with high accuracy [Zhang2020_32936825]. Similarly, ensemble frameworks like DeEPsnap leverage multi-omics data to achieve state-of-the-art performance in essential gene prediction [Zhang2025_40691502]. However, these methods often face challenges related to imbalanced datasets and the interpretability of predictions [Tian2018_30563825].
@@
Recent research has highlighted the importance of integrating network-based features with machine learning models to improve predictive performance [Senthamizhan2021_34630517]. For example, NetGenes, a database of essential genes predicted using interaction network features, demonstrates the utility of network-based approaches in identifying essential genes across diverse bacterial species [Senthamizhan2021_34630517]. Additionally, studies have shown that combining centrality measures with evolutionary properties can enhance the prediction of cancer-associated genes [Safadi2024_38649399].
@@
Despite these advancements, several limitations persist in current approaches. Many methods lack interpretability, making it difficult to understand the biological mechanisms underlying their predictions [Bazaga2020_32612205]. Furthermore, the integration of diverse data types, such as gene expression and epigenetic markers, remains underexplored [Campos2020_32489524]. There is also a need for more robust frameworks that can prioritize therapeutic targets with high confidence and transparency [Zhang2025_40691502].
@@
Emerging trends in bioinformatics emphasize the use of explainable AI techniques, such as SHAP (SHapley Additive exPlanations) [lundberg2017unified], to enhance model interpretability [hsieh2024comprehensive]. These methods provide insights into the contribution of individual features to model predictions, facilitating more informed decision-making in therapeutic target prioritization [mumuni2025explainable].
@@ 2.3 Node Embedding Generation
We applied the Node2Vec [grover2016node2vec] algorithm to generate 128-dimensional vector representations that capture latent network topology features beyond traditional centrality measures. Node2Vec performs biased random walks on the network graph, with walk behavior controlled by two key parameters: the return parameter p and the in-out parameter q. We set both parameters to 1.0 to achieve an unbiased exploration that balances between breadth-first sampling (capturing local neighborhood structure) and depth-first sampling (capturing global network patterns). Each random walk had a length of 80 steps, and we generated 10 walks starting from each node to ensure comprehensive sampling of the network topology around each gene.
@@ 2.4 Gene Essentiality Labels and Target Definition
We used DepMap CRISPR-Cas9 gene knockout screening data as the gold standard for gene essentiality across cancer cell lines [arafeh2025present]. Using these population-scale human gene perturbation datasets enables assessment of essentiality within a physiologically relevant context, consistent with translational model qualification principles. The DepMap project systematically performs genome-wide CRISPR screens across hundreds of cancer cell lines, measuring the fitness effect of knocking out each gene. Essentiality scores represent the median effect across all cell lines, where more negative values indicate greater essentiality for cell survival and proliferation. This pan-cancer approach captures genes that are broadly essential across different cancer types, making our predictions relevant for general therapeutic target discovery rather than being limited to specific cancer subtypes.
@@ 2.6 SHAP-Based Gene Prioritization and Blended Scoring
$\text{SHAP}_{\text{norm}}=\frac{|\text{SHAP}_{\text{sum}}|-|\text{SHAP}_{\text{sum}}|_{\text{min}}}{|\text{SHAP}_{\text{sum}}|_{\text{max}}-|\text{SHAP}_{\text{sum}}|_{\text{min}}}$ $\alpha=0.7\text{ is a weighting parameter.}$
@@ 4.3 Biological Validation and Clinical Relevance
Our framework’s successful prioritization of ribosomal proteins RPS27A, RPS17, and RPS6 as essential genes aligns with mounting evidence for their critical roles in cancer biology and therapeutic potential. RPS27A has emerged as a particularly compelling target, with recent studies demonstrating its prognostic significance in HPV16-positive cervical cancer [wang2021high], where it serves as both a biomarker for patient outcomes and a functionally relevant driver of cancer cell growth. Similarly, RPS17 has been identified as a key component in colorectal cancer stem cell (CRCSC) biology [wu2024integration], where it contributes to the stemness properties that drive tumor metastasis, recurrence, and chemotherapy resistance—making it an attractive therapeutic target for overcoming treatment-resistant cancer phenotypes. RPS6, as a component of the 40S ribosomal subunit, plays dual roles in both ribosomal function and extra-ribosomal signaling, particularly as a downstream effector and surrogate marker of the PI3K/AKT/mTORC1 pathway activation that is prevalent across multiple cancer types [yi2021ribosomal]. The convergent identification of these ribosomal proteins through our network-based approach and their established roles in cancer progression underscores the biological relevance of our predictions and highlights the therapeutic potential of targeting ribosomal dysfunction in cancer. These findings support the growing recognition that ribosomal proteins extend beyond their traditional housekeeping functions to serve as critical regulators of oncogenic processes, making them promising candidates for precision cancer therapy development.
Similarly, the high ranking of MYC confirms our model’s ability to identify oncogenes that are critical for cancer cell survival but challenging to target directly [whitfield2025myc], suggesting that our approach could help identify druggable regulators or downstream effectors of such "undruggable" targets.

### v2

@@ 1 Introduction
The identification of essential genes and therapeutic targets represents a cornerstone of modern biomedical research, with profound implications for understanding fundamental biological processes and developing novel cancer treatments [1]. Essential genes, defined as those critical for cellular survival and reproduction, have been extensively studied across various organisms, from bacteria to humans [2]. Contemporary advances increasingly emphasize human-based experimental and computational models as more predictive and ethically preferable alternatives to traditional animal systems. In cancer research, the prioritization of therapeutic targets is particularly crucial, as it directly influences drug discovery pipelines and treatment efficacy [3]. However, the complexity of biological systems and the vast diversity of human molecular and network data, coupled with population-level variability in genetic and epigenetic susceptibility, presents both a challenge and an opportunity for building integrative human-relevant models. [4].
@@
Protein-protein interaction (PPI) networks have emerged as powerful tools for understanding gene essentiality and identifying potential therapeutic targets [5]. These networks capture the intricate web of molecular interactions within cells, providing insights into cellular functions and disease mechanisms [6]. Centrality measures derived from PPI networks, such as degree, betweenness, and closeness centrality, have been shown to correlate with gene essentiality [7]. Moreover, advancements in network embedding techniques, such as Node2Vec, have enabled the extraction of latent topological features that enhance predictive accuracy [6].
@@
Previous prediction approaches often relied on isolated data modalities or species-agnostic features, lacking the combinatorial integration across multiple in silico modalities that can capture the biological complexity of human systems [8]. For instance, DeepHE, a deep learning-based method, integrates sequence features with PPI network data to predict human essential genes with high accuracy [8]. Similarly, ensemble frameworks like DeEPsnap leverage multi-omics data to achieve state-of-the-art performance in essential gene prediction [5]. However, these methods often face challenges related to imbalanced datasets and the interpretability of predictions [9].
@@
Recent research has highlighted the importance of integrating network-based features with machine learning models to improve predictive performance [10]. For example, NetGenes, a database of essential genes predicted using interaction network features, demonstrates the utility of network-based approaches in identifying essential genes across diverse bacterial species [10]. Additionally, studies have shown that combining centrality measures with evolutionary properties can enhance the prediction of cancer-associated genes [4].
@@
Despite these advancements, several limitations persist in current approaches. Many methods lack interpretability, making it difficult to understand the biological mechanisms underlying their predictions [3]. Furthermore, the integration of diverse data types, such as gene expression and epigenetic markers, remains underexplored [11]. There is also a need for more robust frameworks that can prioritize therapeutic targets with high confidence and transparency [5].
@@
Emerging trends in bioinformatics emphasize the use of explainable AI techniques, such as SHAP (SHapley Additive exPlanations) [12], to enhance model interpretability [13]. These methods provide insights into the contribution of individual features to model predictions, facilitating more informed decision-making in therapeutic target prioritization [14].
@@ 2.3 Node Embedding Generation
We applied the Node2Vec [15] algorithm to generate 128-dimensional vector representations that capture latent network topology features beyond traditional centrality measures. Node2Vec performs biased random walks on the network graph, with walk behavior controlled by two key parameters: the return parameter p and the in-out parameter q. We set both parameters to 1.0 to achieve an unbiased exploration that balances between breadth-first sampling (capturing local neighborhood structure) and depth-first sampling (capturing global network patterns). Each random walk had a length of 80 steps, and we generated 10 walks starting from each node to ensure comprehensive sampling of the network topology around each gene.
@@ 2.4 Gene Essentiality Labels and Target Definition
We used DepMap CRISPR-Cas9 gene knockout screening data as the gold standard for gene essentiality across cancer cell lines [16]. Using these population-scale human gene perturbation datasets enables assessment of essentiality within a physiologically relevant context, consistent with translational model qualification principles. The DepMap project systematically performs genome-wide CRISPR screens across hundreds of cancer cell lines, measuring the fitness effect of knocking out each gene. Essentiality scores represent the median effect across all cell lines, where more negative values indicate greater essentiality for cell survival and proliferation. This pan-cancer approach captures genes that are broadly essential across different cancer types, making our predictions relevant for general therapeutic target discovery rather than being limited to specific cancer subtypes.
@@ 2.6 SHAP-Based Gene Prioritization and Blended Scoring
$\text{SHAP}_{\text{norm}}=\frac{|\text{SHAP}_{\text{sum}}|-|\text{SHAP}_{\text{sum}}|_{\text{min}}}{|\text{SHAP}_{\text{sum}}|_{\text{max}}-|\text{SHAP}_{\text{sum}}|_{\text{min}}}$
$\alpha=0.7\text{ is a weighting parameter.}$
@@ 4.3 Biological Validation and Clinical Relevance
Our framework’s successful prioritization of ribosomal proteins RPS27A, RPS17, and RPS6 as essential genes aligns with mounting evidence for their critical roles in cancer biology and therapeutic potential. RPS27A has emerged as a particularly compelling target, with recent studies demonstrating its prognostic significance in HPV16-positive cervical cancer [17], where it serves as both a biomarker for patient outcomes and a functionally relevant driver of cancer cell growth. Similarly, RPS17 has been identified as a key component in colorectal cancer stem cell (CRCSC) biology [18], where it contributes to the stemness properties that drive tumor metastasis, recurrence, and chemotherapy resistance—making it an attractive therapeutic target for overcoming treatment-resistant cancer phenotypes. RPS6, as a component of the 40S ribosomal subunit, plays dual roles in both ribosomal function and extra-ribosomal signaling, particularly as a downstream effector and surrogate marker of the PI3K/AKT/mTORC1 pathway activation that is prevalent across multiple cancer types [19]. The convergent identification of these ribosomal proteins through our network-based approach and their established roles in cancer progression underscores the biological relevance of our predictions and highlights the therapeutic potential of targeting ribosomal dysfunction in cancer. These findings support the growing recognition that ribosomal proteins extend beyond their traditional housekeeping functions to serve as critical regulators of oncogenic processes, making them promising candidates for precision cancer therapy development.
Similarly, the high ranking of MYC confirms our model’s ability to identify oncogenes that are critical for cancer cell survival but challenging to target directly [20], suggesting that our approach could help identify druggable regulators or downstream effectors of such "undruggable" targets.